Ques 7.
A music store is selling compact disks for $11.50 and $7.50. You buy
12 disks and spend a total of $110. How many compact disks that cost
$11.50 did you buy?

I need these answers to be as straightforward as possible. No shown work is needed. *Thank you for all simply put answers*

Answer by texttutoring(324)

Let x = number of $11.50 discs you bought
let y = number of $7.50 discs you bought

Two equations:
x+y=12 (because there are 12 disks bought in total)
11.50x + 7.50y = 110 (because the total cost is $110)

Use substitution or elimination to solve. I will use substitution.

y = 12-x

11.50x +7.50(12-x) = 110
11.50x + 90 -7.5x = 110
4x = 20
x = 5

So there were five $11.50 disks bought. This means there must have been seven $7.50 disks bought.